Anita can paint the shed in 5 hours. If Kris helps her, they can paint the shed in 3 hours. How long would it take Kris to paint the shed alone?
Answer by stanbon(75887)

Anita can paint the shed in 5 hours. If Kris helps her, they can paint the shed in 3 hours. How long would it take Kris to paint the shed alone?
----
Anita rate = 1/5 job/hr
Together rate = 1/3 job/hr
Kris rate = 1/x job/hr
---------------------------------
Equation:
rate + rate = together rate
1/x + 1/5 = 1/3
----
Multiply thru by 15x to get:
15 + 3x = 5x
2x = 15
x = 7.5 hrs (time for Kris to do the job alone)
=====================================================
Cheers,
Stan H.
======================